Draw a graph showing the effect of a $1 per unit tax on bicycles using our standardconsumer optimization graph. Label the tax revenue and deadweight loss. Assumethe individual has an income of $80 and the price of a bike is $4. Label all axes,curves, and points of interest with as much precision as possible. Note, there will besome things which can’t be assigned a numerical value. Hint: Your Y-Axis shouldbe labeled, “Other Goods.”
Answer #1

Let,

x = quantity of bicycles

y = other goods

p1 = price of bicycle

Price of other goods = $1

M = Income = $80

Budget line,

p1x + y = M = 80

=> 4x + y = 80

After tax on bicycles = t = $1

Budget line,

(p1 + 1)x + y = 80

=> 5x + y = 80

Suppose quantity of bicycle consumed falls from x to x’.

*Please don’t forget to hit the thumbs up button, if you find the answer helpful.

Thank You